# Large Language Model Enabled Multi-Task Physical Layer Network

### v1

@@ V Conclusions
In this paper, we propose a LLM-enabled multi-task physical layer network to unify multiple tasks with a single LLM. Multi-task instruction module, input encoders, as well as output decoder, are elaborately designed to distinguish multiple tasks and adapted the features of different formats of wireless data for the feature of LLM. Simulation results have verified the effectiveness of the proposed method. The proposed LLM framework is promising to perform different tasks according to the requirements with a single model, significantly saving the redundancy and costs of designing LLM for each task separately. Future works can be focused on incorporating more tasks into the network.

### v2

@@ VI Conclusions
In this paper, we propose a LLM-enabled multi-task PHY network to unify multiple tasks with a single LLM. Multi-task instruction module, input encoders, as well as output decoders, are elaborately designed to distinguish multiple tasks and adapted the features of different formats of wireless data for the feature of LLM. Moreover, to reduce the memory requirements of the proposed model, a LoRA fine-tuning-aware quantization method is introduced. Simulation results have verified the effectiveness of the proposed method. The proposed LLM framework is promising to perform different tasks using a single model, significantly saving the redundancy and costs of the practical deployment of LLM. It makes an initial attempt to provide a more adaptable, comprehensive, and intelligent PHY network with the aid of LLMs. Future works can be focused on incorporating more tasks into the network. Besides, the LLM pruning and knowledge distillation techniques can also be employed to reduce the computational and storage overhead of LLM-based method.